OpenStudy (anonymous):

what is y 3/5 as a radical

OpenStudy (anonymous):

radical means root \[\sqrt{x}\]

OpenStudy (anonymous):

\[\sqrt[5]{y^3}\]

OpenStudy (anonymous):

Think of it as the bottom or denominator goes outside the root and the numerator as the power of what is inside the root

OpenStudy (mathmale):

Sarah, I know it can be a challenge to type some kinds of math notation from an ordinary keyboard, but accuracy in presentation is important. Your y 3/5 would be better presented as y^(3/5) or \[y^{3/5}\]or 

OpenStudy (mathmale):

Mateaus has given you the best possible way in which to express your post properly in radical form:\[\sqrt[5]{y^3}\]

OpenStudy (anonymous):

or \[(\sqrt[5]{y})^{3}\] :P

OpenStudy (mathmale):

That's definitely an acceptable alternative.